OpenStudy (anonymous):

Write an equation for a line perpendicular to y=5x+1 and passing through the point (-5,-2) PLease

hartnn (hartnn):

can u find slope of y=5x+1 ?

OpenStudy (anonymous):

yea 5/1

hartnn (hartnn):

good, now slope of perpendicular line will be reciprocal inverse. like if slope was m, slope of perpendicular line = -1/m here , slope of perpendicular line will be ... ?

OpenStudy (anonymous):

is it y=-1/5x-23??

OpenStudy (anonymous):

+

hartnn (hartnn):

-23?? (1/5x is correct) y= -1/5 x +b

OpenStudy (anonymous):

+23

hartnn (hartnn):

y= -1/5 x +b put x = -5, y= -2 and find b and b does not =23

OpenStudy (anonymous):

-1??

hartnn (hartnn):

lets check. -2 = (-5/-5) +b -2 =1+b b= -2-1 =... ?

OpenStudy (anonymous):

oh nvm it -3

OpenStudy (anonymous):

-1/5x-3

hartnn (hartnn):

yes. y= -1/5 x -3

OpenStudy (anonymous):

omg thank u